# From Age Estimation to Age-Invariant Face Recognition: Generalized Age Feature Extraction Using Order-Enhanced Contrastive Learning

### v1

@@ III Order-Enhanced Contrastive Learning
In this section, we provide a detailed explanation of the proposed OrdCon, beginning with the formulation of age feature extraction. We then explain the contrastive learning objectives based on order learning and metric learning, followed by a discussion on how the method can be applied to age estimation and AIFR. Fig. 2 shows the workflow of the proposed method.
@@ III-A Problem Formulation
After feature extraction, age estimation can be formulated as a regression problem, where the goal is to predict the chronological age $\hat{y}_{age}^{(i)}\in\mathbb{R}$ from a given facial image $x^{(i)}$. The model is trained to minimize the error between the predicted age $\hat{y}_{age}^{(i)}$ and the ground-truth age $y_{age}^{(i)}$.
In contrast, the objective of AIFR is to preserve identity across different ages. Given a set of facial images $\{x^{(j)}\}_{j=1}^{M}$ of the same individual taken at different ages, the model should map these images to a shared identity feature space $\mathcal{Z}_{id}$, where identity remains invariant to age-related changes. Let $z^{(j)}_{id}=f_{id}(x^{(j)})$ represent the identity feature vector extracted from image $x^{(j)}$, where $f_{id}$ is the learned mapping function. To ensure identity preservation, the model must minimize the distance among the identity features $\{z^{(j)}_{id}\}_{j=1}^{M}$.
@@ III-B1 Feature Direction Contrast through Order Learning
Conversely, if $y_{age}^{(i)}>y_{age}^{(j)}$, the age relationship is regressive, denoted by $x^{(i)}\succ x^{(j)}$, meaning $x^{(i)}$ represents an older face than $x^{(j)}$. In this case, the feature space should reflect the correct temporal ordering by placing $z_{age}^{(i)}$ ahead of $z_{age}^{(j)}$. Lastly, the order learning objective ignores pairs where $y_{age}^{(i)}=y_{age}^{(j)}$, as there is no order information to learn. Instead, the metric learning objective minimizes the distance between age features for these images to ensure that features of the same age are closely grouped, reducing intra-class variance.
@@
where $v_{f}^{(i,j)}$ represents the forward reference direction, and $v_{b}^{(i,j)}$ represents the backward reference direction. The order learning objective is to maximize the similarity between $v_{f}^{(i,j)}$ and $v^{(i,j)}_{d}$, aligning the direction vector with the direction of age progression, while minimizing the similarity between $v_{b}^{(i,j)}$ and $v^{(i,j)}_{d}$. This is achieved by optimizing the following loss function:
@@
$\tau$ is a scaling hyper-parameter that controls the sensitivity to differences between positive and negative pairs. $\mathbbm{1}_{(\cdot)}$ is an indicator that equals to 1 if the condition in the parenthesis is true and 0 otherwise.
@@
The order learning objective, in this case, aims to minimize the similarity between $v_{f}^{(i,j)}$ and $v^{(i,j)}_{d}$ while maximizing the similarity between $v_{b}^{(i,j)}$ and $v^{(i,j)}_{d}$ by optimizing the regressive counterpart:
@@
The order learning objective is defined as the sum of both terms:
@@
It is important to note that no weighted hyper-parameters are used, as both progressive and regressive relationships are equally important to learn.
@@ III-B2 Soft Proxy Matching Loss
While order learning ensures that the model captures the natural progression of aging by aligning age features with reference directions, it does not directly address intra-class variance or inter-class separation. To address this, we use metric learning with proxies to cluster similar age features together while ensuring that features from different ages remain distinct.
As previously discussed, each distinct age $a\in\mathcal{A}$ is associated with a proxy $c_{a}\in\mathbb{R}^{d_{age}}$. In proxy-based metric learning, these proxies act as representative centers for each age in the feature space, pulling features of the same age closer to their proxy while pushing them away from proxies representing other ages.
Unlike the loss functions proposed in [24] and [69], which use Euclidean distances, we propose a proxy matching mechanism with cosine similarities to better capture the direction of age features, which is crucial for modeling age progression. For each given $x^{(i)}$, the proxy matching loss function is formulated as:
$\mathcal{L}_{pm}(x^{(i)},\mathcal{C}^{(i)})=\frac{exp(sim(x^{(i)},c_{y_{age}^{(i)}})/{\tau})}{\sum_{\begin{subarray}{c}z\in\mathcal{Z}\end{subarray}}exp(sim(x^{(i)},c_{z})/{\tau})},$ (13)
where $sim(\cdot)$ represents the cosine similarity. $\mathcal{Z}$ is a set of proxies such that $c_{y_{age}^{(i)}}\cup\mathcal{Z}=\mathcal{A}$ and $c_{y_{age}^{(i)}}\cap\mathcal{Z}=\varnothing$.
@@
While cosine similarity effectively captures the direction of age features, it does not account for the magnitude of different age differences. For instance, the distance between an age feature and a proxy with lower similarity should be increased more than that between those with higher similarity. To address this, we propose the soft proxy matching loss, which assigns a weight term to each negative pair in the denominator of the loss function in Eq. (13). The soft proxy matching loss is formulated as:
$\mathcal{L}_{spm}(x^{(i)},\mathcal{C}^{(i)})=\frac{exp(sim(x^{(i)},c_{y_{age}^{(i)}})/{\tau})}{\sum_{\begin{subarray}{c}z\in\mathcal{Z}\end{subarray}}w(y_{age}^{(i)},z)exp(sim(x^{(i)},c_{z})/{\tau})},$ (14)
@@
The metric learning objective is defined as the aggregation of soft proxy matching losses for all input samples:
$\mathcal{L}_{metric}=-\frac{1}{N}\sum_{\begin{subarray}{c}i=1\end{subarray}}^{N}\mathcal{L}_{spm}(x^{(i)},\mathcal{C}^{(i)}),$ (16)
The overall contrastive loss is a weighted sum of the order learning objective and the metric learning objective, formulated as:
$\mathcal{L}_{contrast}=\mathcal{L}_{order}+\lambda\mathcal{L}_{metric},$ (17)
@@ III-C Application to Age Estimation and AIFR
The proposed OrdCon is versatile and can be applied to two key age-related tasks: age estimation and AIFR. As shown in Fig.2 OrdCon learns age features that can be directly used for age estimation. Once age progression is modeled during the contrastive learning stage, a regression loss, such as L1 loss, can be used to map the learned features to their corresponding labels. For age estimation, the L1 loss is formulated as:
@@
Unlike [25], where a GAN [70] generates an additional sample within a random age group as part of the data augmentation process, OrdCon generates samples for every age group except the one represented by the input face to learn the identity of the same individual across different ages. To achieve this, a condition (the age group label) is fed to the GAN to guide the synthesis process. We use the same GAN model as in [70] due to its superior capability in identity preservation, which is crucial for face recognition with synthesized samples [71].
@@ IV-A Datasets
We evaluate OrdCon on three commonly used age-oriented benchmark datasets, which include age and identity labels suitable for both age estimation and AIFR tasks: the MORPH II dataset [35], the FG-NET dataset [73], and the AgeDB dataset [74]. The statistics for these datasets are listed in Table I. For both tasks, we conduct both homogeneous-dataset and cross-dataset experiments to demonstrate the robustness and versatility of our method. Some state-of-the-art methods use the IMDB-WIKI dataset [17] for pre-training; however, we do not use it, as we observed no significant performance improvement with our proposed method. When making comparisons, we explicitly indicate whether a method uses the IMDB-WIKI dataset for pre-training.
@@
The AgeDB dataset contains 16,516 facial images of 570 celebrities, with ages ranging from 0 to 101 years. The images also exhibit variations in PIE. In our experiments, we use only the AgeDB dataset for celebrities, as other celebrity datasets have noisy labels and overlapping identities, making them unsuitable for cross-dataset evaluations. Notably, we use the AgeDB dataset only for cross-dataset evaluations and ablation studies, as performance on this dataset is less frequently reported compared to the MORPH II and FG-NET datasets.
@@ IV-B1 Data Pre-processing
We use the open-source computer vision library dlib [75] for image preprocessing. Initially, 68 facial landmarks are detected in each image, which is then cropped based on eye positions and resized to $128\times{128}$ pixels.
@@ IV-B2 Data Partition
For age estimation experiments with the MORPH II dataset, three commonly used experimental settings are applied. In the first setting, referred to as Setting I, following previous works [43, 46, 45, 44], the dataset is randomly split into two subsets, with 80% used for training and 20% for testing, ensuring no identity overlap. We generate 10 different partitions (same ratio but different splits) and report the mean values. In the second setting (Setting II), to address the imbalance in race distribution, the dataset is divided into three subsets: S1, S2, and S3, maintaining a 1:1 ratio between Black and White subjects and a 3:1 ratio between Male and Female subjects [76, 42, 40, 34]. The goal is to select all females while maximizing the overall size of the subset. We train the model on S1 and test on S2+S3, then train on S2 and test on S1+S3, and report the average result. Lastly, in Setting III, to reduce the variance caused by race imbalance, 5,492 images of White individuals are selected [17, 37]. These images are randomly split into training (80%) and testing (20%), and final results are obtained using 5-fold cross-validation to further minimize data distribution variance.
@@ IV-B3 Implementation Details
For fair comparison with other methods, we use ResNet-50 [82] as the feature extractor for both age estimation and AIFR experiments. Fully connected layers of size 2048 are used to generate age and identity features for contrastive losses. During fine-tuning, an additional fully connected layer is added, with size 1 for age estimation or a size equal to the number of identities in the dataset for identity recognition. The batch size is set to 4,096 during pre-training and 512 during fine-tuning. For contrastive pre-training, we run 400 epochs for age estimation and 1,000 epochs for AIFR. After epoch 500 in AIFR, we gradually adjust $\lambda_{grl}$ using the following equation:
@@
where $t$ represents the training epoch, and $\gamma$ controls the growth rate of $\lambda_{grl}$, set to 10 for all AIFR experiments. $\lambda$ is set to 0.8 for all experiments. Additionally, we use the LARS optimizer for multi-GPU training during pre-training and Stochastic Gradient Descent (SGD) for fine-tuning.

### v2

@@ III Order-Enhanced Contrastive Learning
In this section, we provide a detailed explanation of the proposed OrdCon, beginning with the formulation of the age feature extraction. We then explain the contrastive learning objectives, followed by a discussion on how the method can be applied to age estimation and AIFR. Fig. 2 shows the workflow of the proposed method.
@@ III-A Problem Formulation
After feature extraction, age estimation can be formulated as a regression problem, where the goal is to predict the chronological age $\hat{y}_{age}^{(i)}\in\mathbb{R}$ from a given facial image $x^{(i)}$. The model is trained to minimize the error between the predicted age $\hat{y}_{age}^{(i)}$ and the ground-truth age $y_{age}^{(i)}$, and the learned features should be positioned in the feature space align with the direction of the age progression.
In contrast, AIFR aims to preserve identity across different ages. Given a set of facial images $\{x^{(j)}\}_{j=1}^{M}$ of the same individual taken at different ages, the model should map these images to a shared identity feature space $\mathcal{Z}_{id}$, where identity remains invariant to age-related changes. Let $z^{(j)}_{id}=f_{id}(x^{(j)})$ represent the identity feature vector extracted from image $x^{(j)}$, where $f_{id}$ is the learned mapping function. To ensure identity preservation, the model must minimize the distance among the identity features $\{z^{(j)}_{id}\}_{j=1}^{M}$.
@@ III-B1 Feature Direction Contrast through Order Learning
Conversely, if $y_{age}^{(i)}>y_{age}^{(j)}$, the age relationship is regressive, denoted by $x^{(i)}\succ x^{(j)}$, meaning $x^{(i)}$ represents an older face than $x^{(j)}$. In this case, the feature space should reflect the correct temporal ordering by placing $z_{age}^{(i)}$ ahead of $z_{age}^{(j)}$. Lastly, the order learning objective ignores pairs where $y_{age}^{(i)}=y_{age}^{(j)}$, as there is no order information to learn. Instead, the proxy-based contrastive objective minimizes the distance between age features for these images, ensuring that features of the same age are closely grouped.
@@
where $v_{f}^{(i,j)}$ represents the forward reference direction, and $v_{b}^{(i,j)}$ represents the backward reference direction. The order learning-based contrastive objective is to maximize the similarity between $v_{f}^{(i,j)}$ and $v^{(i,j)}_{d}$, aligning the direction vector with the direction of age progression, while minimizing the similarity between $v_{b}^{(i,j)}$ and $v^{(i,j)}_{d}$. This is achieved by optimizing the following loss function:
@@
$\tau$ is a scaling hyperparameter that controls the sensitivity to differences between positive and negative pairs. $\mathbbm{1}_{(\cdot)}$ is an indicator that equals one if the condition in the parenthesis is true and zero otherwise.
@@
The order learning-based contrastive objective, in this case, aims to minimize the similarity between $v_{f}^{(i,j)}$ and $v^{(i,j)}_{d}$ while maximizing the similarity between $v_{b}^{(i,j)}$ and $v^{(i,j)}_{d}$ by optimizing the regressive counterpart:
@@
The aggregated order learning-based contrastive objective is defined as the sum of both terms:
@@
It is important to note that no weighted hyperparameters are used, as both progressive and regressive relationships are equally important to learn.
By incorporating order learning, OrdCon learns not only that samples are different, but also how they differ in an ordinal sense (e.g., elder/younger). This imposes a meaningful structure on the learned feature space, aligning it with the natural progression of aging. Such a structured representation is inherently more generalizable than simply mapping features to labels without considering their underlying relationships.
@@ III-B2 Soft Proxy Matching Loss
While order learning-based objective ensures that the model captures the natural progression of aging by aligning age features with reference directions, it does not directly address intra-class variance or inter-class separation, which are keys to good generalizability [19]. To address this, we propose a new contrastive objective with proxies to cluster similar age features together while ensuring that features from different ages remain distinct.
As previously discussed, each distinct age $a\in\mathcal{A}$ is associated with a proxy $c_{a}\in\mathbb{R}^{d_{age}}$. In proxy-based representation learning, these proxies act as representative centers for each age in the feature space, pulling features of the same age closer to their proxy while pushing them away from proxies representing other ages.
Unlike the loss functions proposed in [56] and [57], which utilize Euclidean distances, we propose a proxy matching mechanism based on cosine similarities to better capture the direction of age features, which is crucial for modeling age progression. For each given $x^{(i)}$, the proxy matching loss function is formulated as:
$\mathcal{L}_{pm}(x^{(i)},\mathcal{C})=\frac{exp(sim(x^{(i)},c_{y_{age}^{(i)}})/{\tau})}{\sum_{\begin{subarray}{c}z\in\mathcal{Z}\end{subarray}}exp(sim(x^{(i)},c_{z})/{\tau})},$ (13)
where $sim(\cdot)$ represents the cosine similarity and $c_{y_{age}^{(i)}}$ indicates the corresponding proxy $x^{(i)}$ should approach to. $\mathcal{Z}$ is a set of proxies such that $c_{y_{age}^{(i)}}\cup\mathcal{Z}=\mathcal{A}$ and $c_{y_{age}^{(i)}}\cap\mathcal{Z}=\varnothing$.
@@
While cosine similarity effectively captures the direction of age features, it does not account for the magnitude of different age differences. From a generalization perspective, the distance between an age feature and a proxy with lower similarity should be increased more than that between those with higher similarity. To address this, we propose the soft proxy matching loss, which assigns a weight term to each negative pair in the denominator of the loss function in Eq. (13). The soft proxy matching loss is formulated as:
$\mathcal{L}_{spm}(x^{(i)},\mathcal{C})=\frac{exp(sim(x^{(i)},c_{y_{age}^{(i)}})/{\tau})}{\sum_{\begin{subarray}{c}z\in\mathcal{Z}\end{subarray}}w(y_{age}^{(i)},z)exp(sim(x^{(i)},c_{z})/{\tau})},$ (14)
@@
The proxy-based objective is defined as the aggregation of soft proxy matching losses for all input samples:
$\mathcal{L}_{proxy}=-\frac{1}{N}\sum_{\begin{subarray}{c}i=1\end{subarray}}^{N}\mathcal{L}_{spm}(x^{(i)},\mathcal{C}).$ (16)
This new loss function ensures features of the same age cluster are tightly clustered while pushing features of different ages apart proportionally to their age difference. This creates a more discriminative and well-organized feature space, which is crucial for generalization, especially in cross-dataset scenarios where data distributions might differ.
The overall contrastive loss is a weighted sum of the order learning objective and the proxy matching objective, formulated as:
$\mathcal{L}_{contrast}=\mathcal{L}_{order}+\lambda\mathcal{L}_{proxy},$ (17)
@@ III-C Application to Age Estimation and AIFR
The proposed OrdCon is a new contrastive learning framework, and we apply it to two key age-related tasks: age estimation and AIFR. As shown in Fig.2, OrdCon learns age features that can be directly used for age estimation. Once age progression is modeled during the contrastive learning stage, a regression loss, such as L1 loss, can be used to map the learned features to their corresponding labels. For age estimation, the L1 loss is formulated as:
@@
Unlike [26], where a generative adversarial network (GAN) [58] synthesizes an additional sample within a random age group as part of the data augmentation process, OrdCon leverages samples for every age group except the one represented by the input face to learn the identity of the same individual across different ages. To achieve this, a condition (the age group label) is fed to the GAN to guide the synthesis process. We use the same GAN model as in [58] due to its superior capability in identity preservation, which is crucial for face recognition with synthesized samples [59].
@@
III-D Generalizability Evaluation Matrices
Based on the objectives of age estimation and AIFR mentioned in Section III.A, we adapt two matrices, alignment of positive samples and divergence of class centers, from [19] to test the generalizability of OrdCon. Alignment of positive samples measures the average squared Euclidean distance between feature vectors of two augmented views, and divergence of class centers is defined by the average dot product between centers of every pair of different classes. Since they are designed for classification problems with categorical data, they can not be used directly on tasks involving ordinal data. To this end, we evaluate the alignment of positive samples on AIFR and modify the divergence of class centers to make it applicable to age estimation. The alignment of positive samples for AIFR is formulated as:
$\mathcal{A}lign=\frac{1}{N}\sum_{\begin{subarray}{c}i=1\end{subarray}}^{N}\|f_{id}(A_{1}(x^{(i)}))-f_{id}(A_{2}(x^{(i)}))\|^{2},$ (23)
where $A_{1}(x^{(i)})$ and $A_{2}(x^{(i)})$ are two augmented samples of $x^{(i)}$. A small alignment means positive pairs are closer in the feature space, which is desirable.
Regarding divergence, instead of calculating the dot product between class centers, we take the dot product between direction vectors and the overall age progression direction. The overall age progression direction in this case is defined by the direction vector of the youngest and eldest class centers. We call this new matrix the divergence of direction vectors, and it is formulated as:
$\mathcal{D}iv=\frac{1}{|\mathcal{P}|}\sum_{(i,j)\in\mathcal{P}}\left((v^{(i,j)})^{\top}\cdot v_{overall}\right),$ (24)
where $\mathcal{P}$ denotes the set of all progressive pairs, $v_{overall}$ is the overall age progression direction and is defined as:
$v_{overall}=\frac{c_{eldest}-c_{youngest}}{||c_{eldest}-c_{youngest}||}.$ (25)
$c_{eldest}$ and $c_{youngest}$ are the class centers of the eldest and youngest ages in the dataset. Instead of the smaller the better when measuring the divergence of class centers, a divergence of direction vectors close to 1 is ideal, which means the directions of progressive pairs are in the same direction as the overall aging trend in the feature space.
@@ IV-A Datasets
We evaluate OrdCon on three commonly used age-oriented benchmark datasets, which include age and identity labels suitable for both age estimation and AIFR tasks: the MORPH II dataset [40], the FG-NET dataset [61], and the AgeDB dataset [62]. The statistics for these datasets are listed in Table I. For both tasks, we conduct both homogeneous-dataset and cross-dataset experiments to demonstrate the robustness and generalizability of our method. Some state-of-the-art methods use the IMDB-WIKI dataset [14] for pre-training; however, we do not use it, as we observed no significant performance improvement with our proposed framework. When making comparisons, we explicitly indicate whether a method uses the IMDB-WIKI dataset for pre-training.
@@
The AgeDB dataset contains 16,516 facial images of 570 celebrities, with ages ranging from 0 to 101 years. The images also exhibit variations in PIE. In our experiments, we use only the AgeDB dataset for celebrities, as other celebrity datasets have noisy labels and overlapping identities, making them unsuitable for cross-dataset evaluations. Notably, we utilize the AgeDB dataset only for cross-dataset evaluations and ablation studies, as its performance is less frequently reported compared to that of the MORPH II and FG-NET datasets.
@@ IV-B1 Data Pre-processing
We use the open-source computer vision library dlib [63] for image preprocessing. Initially, 68 facial landmarks are detected in each image, which is then cropped based on eye positions and resized to $\times{128}$ 128 pixels.
@@ IV-B2 Data Partition
For age estimation experiments with the MORPH II dataset, three commonly used experimental settings are applied. In the first setting, referred to as Setting I, following previous works [64, 45, 44, 43], the dataset is randomly split into two subsets, with 80% used for training and 20% for testing, ensuring no identity overlap. We generate 10 different partitions (with the same ratio but different splits) and report the mean values. In the second setting (Setting II), to address the imbalance in race distribution, the dataset is divided into three subsets: S1, S2, and S3, maintaining a 1:1 ratio between Black and White subjects and a 3:1 ratio between Male and Female subjects [65, 66, 67, 68]. The goal is to select all females while maximizing the overall size of the subset. We train the model on S1 and test it on S2+S3, then train it on S2 and test it on S1+S3, and report the average result. Lastly, in Setting III, to reduce the variance caused by race imbalance, 5,492 images of White individuals are selected [14, 42]. These images are randomly split into training (80%) and testing (20%) sets, and final results are obtained using 5-fold cross-validation to further minimize data distribution variance.
@@ IV-B3 Implementation Details
For fair comparison with other methods, we use ResNet-50 [75] as the feature extractor for both age estimation and AIFR experiments. Fully connected layers of size 2048 are used to generate age and identity features for contrastive objectives. During fine-tuning, an additional fully connected layer is added, with size 1 for age estimation or a size equal to the number of identities in the dataset for identity recognition. The batch size is set to 4,096 during pre-training and 512 during fine-tuning. For contrastive pre-training, we run 400 epochs for age estimation and 1,000 epochs for AIFR. After epoch 500 in AIFR, we gradually adjust $\lambda_{grl}$ using the following equation:
@@
where $t$ represents the training epoch, and $\gamma$ controls the growth rate of $\lambda_{grl}$, and is empirically set to 10 for all AIFR experiments. $\lambda$ in Eq. (17) is empirically set to 0.8 for all experiments. Additionally, we use the LARS optimizer for multi-GPU training during pre-training and Stochastic Gradient Descent (SGD) for fine-tuning.
@@
IV-C Age Estimation Results
IV-C1 Homogeneous-Dataset Results
The homogeneous-dataset results for the MORPH II and FG-NET datasets are presented in Tables II and III, respectively. Note that in both tables, * indicates results produced by using the official code release. In Table II, results for all three settings are reported. As shown, OrdCon achieves either state-of-the-art or comparable results across all settings. Despite not using the IMDB-WIKI dataset for pretraining, our method outperforms those that do, highlighting the importance of modeling age progression between samples, whereas most methods only learn a direct mapping between features and labels. MCGRL [80], which is also based on contrastive learning, achieves the lowest MAE in two of the three settings. One reason could be that it primarily focuses on learning the facial structure within the dataset, thereby gaining an advantage in homogeneous-dataset experiments. Notably, our method outperforms RNC [25], which also employs contrastive learning to rank samples in order. By incorporating the soft proxy matching loss, our method learns features that are positioned closer to the center of each age, resulting in a more compact feature distribution that aligns with natural age progression. Fig. 4 illustrates how the learned age feature space on the MORPH II dataset evolves during training. It also highlights the dataset’s imbalance, with the majority of younger faces represented by the thicker dark green end and the minority of older faces represented by the lighter green end.
In Table III, the homogeneous-dataset results on the FG-NET dataset show a similar trend to those on the MORPH II dataset. However, due to the small size of the FG-NET dataset, methods pretrained with the IMDB-WIKI dataset achieve relatively lower MAE compared to those trained from scratch. For OrdCon, we found that it does not benefit from pretraining using IMDB-WIKI, as it could not effectively define progressive and regressive pairs based on the noisy labels in that dataset. Despite this, OrdCon achieves comparable results to those of other methods without requiring additional pretraining.